# supplemental_arxiv

### v1

@@ 5 CDC flu forecasts analysis
In this section, the QGP model is used for QM of quantile forecasts targeting US flu hospitalizations. Beginning in 2013, the CDC began hosting a yearly forecast competition of the influenza outbreak in the US. This competition is known as FluSight. The flu epidemic typically begins in the fall and ends in late spring the following year, and the forecast competition lasts around 30 weeks starting in October and ending in May. FluSight involves several academic and industry research teams who each independently develop forecasts every week for predicting certain flu targets for future weeks Biggerstaff et al. (2016). During the 2022-23 and 2023-24 seasons, the forecast targets were the 1, 2, 3, and 4-week ahead hospitalizations as reported by Health and Human Services (HHS) for the 50 US states, Puerto Rico, the District of Columbia, and the nation as a whole. The data for weekly hospitalizations of flu patients may be found at HealthData.gov (2024). The official guidelines of the most 2023-24 FluSight and all submitted quantile forecasts are publicly available at Github (2024) Mathis et al. (2024).
Participating teams were free to create forecasts however they pleased, but forecasts for each target were required to consist of 23 quantiles for the probability levels $\boldsymbol{p}=(0.01,0.025,0.5,0.1,...,0.95,0.975,0.99)$, which were given by FluSight. Figure 4 shows examples of quantile forecasts of the same target from 12 participating teams. Each plot shows the log forecast for hospitalizations in the US for the week of January 13, 2024, and it is clear that there are major distributional differences between forecasts. We denote a quantile forecast for flu hospitalizations as $\hat{Q}^{(H)}(\boldsymbol{p})=(\hat{Q}^{(H)}(0.01),...,\hat{Q}^{(H)}(0.99))$. The general quantile forecast representation format allowed for forecasts to be compared by the same metric, the weighted interval score (WIS), and to be easily combined into a multi-model ensemble forecast Mathis et al. (2024). However, under the quantile representation, the tools for scoring forecasts and for building ensemble models are limited as many of the existing tools for scoring forecasts and constructing ensembles require CDFs or PDFs Wadsworth et al. (2023); Ranjan and Gneiting (2010). Because of these limitations, it may be desirable to approximate continuous distributions from the quantile forecasts to allow for more flexibility in scoring or ensemble building. In this section, we fit the QGP model for QM to every forecast of hospitalizations submitted to the FluSight during the 2023-24 season, and an analysis is made to compare the scoring of the quantile forecasts to QGP estimated forecasts using proper scoring rules.
@@
Before fitting a QGP model to a quantile forecast we transformed the quantiles $\hat{Q}^{(H)}(\boldsymbol{p})$ to $\text{log}(\hat{Q}^{(H)}(\boldsymbol{p})+1)$ so that forecasts for all states were on a similar scale. The model fit was the same as the QGP fit in section 4.2, including the same prior distribution assignments. According to the official forecast competition rules, no forecast could include quantiles that were less than 0. As a result, there were forecasts with one or many quantiles equal to 0. When fitting the QGP model, these instances of 0 values were removed so that some forecasts had $K<23$ quantiles. Between 39 competing forecast teams, 53 locations, 29 forecast dates, and 4 horizons, there were 180,312 quantile forecasts to which the QGP was fit. For each forecast, the WIS was calculated for the quantile forecast and the continuous ranked probability score (CRPS) was calculated over the predictive distribution of the fit QGP model.
The WIS and CRPS are both proper scoring rules, which is a class of scoring rule defined so as to keep a forecaster honest in their forecasts Gneiting and Raftery (2007); Gneiting and Katzfuss (2014). The definition for the WIS is in (19) and is the same as that in Bracher et al. (2021). The WIS consists of the sum of multiple intervals scores (IS), the definition of which is in (20). Here $\alpha$ is the nominal level of an interval with $l$ and $r$ being the respective lower and upper bounds of the interval. $R$ is the number of intervals in the quantile forecast, $\alpha_{r}$ is the nominal level for the $r^{th}$ interval, and $y^{*}$ is the observed value which one is attempting to forecast.
@@
The CRPS is a widely used scoring rule for forecasting and is a function of the CDF of a continuous distribution thus making it unavailable for scoring quantile functions. The CRPS is defined in (21) where $F$ is the CDF of a forecast and $y^{*}$ is the observed event one attempts to forecast. The definition is the same as in Gneiting and Katzfuss (2014).
[…]
@@
Fitting continuous distributions to the quantile forecasts would allow for forecast comparison using the CRPS. The CRPS assesses a forecast across an entire distribution, including the tails, which the WIS is unable to do, giving more reason why one may want to perform QM on a quantile forecast. We fit the QGP model in (12) to the forecast competition forecasts from the 2023-24 season and compare the results of scoring the given quantile forecasts by the WIS and the CRPS calculated from the posterior predictive of the QGP model. Posterior predictive samples from the QM forecasts allow for approximate calculation of the CRPS. To calculate the WIS, we use the evalcast R package McDonald et al. (2023), and to calculate the CRPS we use the scoringutils package Jordan et al. (2019).
Figures 5 and 6 show results of comparing the WIS from quantile forecasts and the CRPS from the QGP fitted forecasts. Figure 5 has in the $x$-axis the WIS and CRPS is in the $y$-axis. Each point is for one forecast for any participating team, season week, and state. The plot is faceted by horizon. Clearly the relationship between the quantile forecasts and the QGP forecasts are very close with high correlation near 0.9 for each horizon.
@@
Figure 5: Scatterplots of WIS and CRPS values for forecasts from 39 different teams for 2023-24 CDC flu forecast competition. The plots are faceted by forecast horizon. Each point represents scores for a flu forecast for one week during the season, one state, and one competing team. Points are transparent to show where more scores tended to be. Overall linear correlation is also given in the corner of each plot.
Figure 6 shows the ranking of teams under the quantile forecasts’ WIS vs. the QGP forecasts’ CRPS. Here we say a team is considered the best performing by having the lowest score averaged across all season weeks, states, and horizons. The $x$-axis indicates where one of the 39 teams ranks by WIS and the $y$-axis where they rank by CRPS. If a block is on the diagonal line, the team’s ranking was the same under the WIS as under the CRPS. The darkness of the block indicates how correlated that team’s WIS is to the CRPS over the whole season. Of the 39 teams represented, 16 have the same ranking for both scores, and for over half of all teams the ranking difference between WIS and CRPS is within 3 positions. For most teams the linear correlation between WIS and CRPS is very high with only five teams having a correlation below 0.85.
Figure 6: Blocks showing the overall season ranking of 39 competing forecast teams for WIS ($x$-axis) and CRPS ($y$-axis). Blocks on the diagonal line have the same ranking under both scores. Blocks are shaded to show correlation between WIS and CRPS for all forecasts submitted by that team.
The results in this section show the close relationship between the quantile forecasts of the CDC flu forecast competition and the more complete forecast distributions estimated via QGP. The implications of these results include being able to score forecasts according to scoring rules used for continuous distributions, including more problem specific rules, and being able to combine forecasts into an ensemble according to sums of CDFs or PDFs. Increased flexibility in scoring and ensemble construction could lead to improved decision making as problem specific evaluation becomes more accessible.
@@ 6 Conclusion
In this paper we consider the situation where quantiles with known probability levels are given as data or is given to represent a distribution rather than raw data or a fully defined distribution. We review basic properties of QFs, sample quantiles, and resulting CLTs which motivate a new model, the QGP model, for matching a set of quantiles to a continuous distribution. The QGP model is assessed for parameter estimation and inference, distribution approximation, and it is compared with other QM methods found in the literature. The simulation studies show that the QGP model does well in estimating parameters and allows for accurate measurement of uncertainty. It provides a way to perform QM for distributions defined by CDFs as well as quantile distributions where the CDF is difficult to evaluate. An application of the QGP for QM of quantile forecasts from the 2023-24 FluSight project shows how QGP matched distributions are closely related to the original quantile forecasts in terms forecast performance among competing models. QGP forecasts, however, may be scored using a variety of scoring rules unavailable to quantile forecasts, and independent forecasts may be combined using methods made only for combining distributions by CDF or PDF functions. QGP has already been applied in chapter 4 of Wadsworth (2024) who used the QGP predictive distributions fit in section 5 to construct ensemble forecasts.
Approximating continuous distributions given only a set of estimated quantiles is not a new idea, but QGP provides a method of doing so while also accurately accounting for asymptotic uncertainty in estimated or predicted quantiles. The CLTs in this paper are limited to the case where the quantiles are estimated given a distribution sample, however similar results have been shown for the quantile regression case where quantiles are estimated in the presence of covariates Kocherginsky et al. (2005); Koenker and Bassett Jr (1978). These results could lead to additional QM modeling using the QGP model where covariates are given, and this may be more akin to the QM done by Sgouropoulos et al. (2015). Additional research for the QGP model may be in how the required distribution function is selected. Where a true model is unknown, we elected to model a distribution as a normal mixture distribution. Of course this has its limits, and using other non parametric functions, similar to Gasthaus et al. (2019), may provide needed flexibility. As long as the selected function is continuous and once differentiable, the theory herein applies.
The application in this paper suggests QM being useful for forecast hubs. Gerding et al. (2023) found QM a necessity in order to evaluate forecasts by a new scoring rule with COVID-19 specific applications. The extra work that comes with QM may or may not be worth the effort for any given forecast hub, but in some contexts it is something worth considering.
7 Acknowledgments
@@
The figures herein are supplementary to those in the main manuscript. Figure 7 shows examples of marginal posterior densities for QM of quantiles for a normal distribution from the simulation study in section 4.1 in the main manuscript. The parameter uncertainty of the QGP and ORD models are unsurprisingly similar in most cases, and where they differ it is hard to say that one is estimating the true parameter better than the other. The IND model on the other hand has tighter densities, and often the bulk of the posterior is far from the true parameter.
Figure 7: Density plots of posterior distribution samples for normal parameters by QM for QGP, ORD, and IND models. QM was done on estimated quantiles from a normal distribution with mean 4 and standard deviation 3.5. The posterior densities are for $\mu$ (top left), $\sigma$ (top right), and sample size $n$ (bottom). Plots are faceted by the sample size $n\in\{50,150,500,1{,}000\}$ ($y$-axis) and number of quantiles $K\in\{7,15,23\}$ ($x$-axis). Vertical lines (black) show the value of the true parameter.
Figures 8 and 9 are from a simulation study similar to that of section 4.1 except that QM was performed on quantiles estimated from draws from an exponential distribution with parameter $\lambda=4$ rather than a normal distribution. The results in the study for the exponential distribution were similar to those for the normal distribution.
Figure 10 was referred to in section 4.1.2 of the main manuscript, and figures 11, 12, and 13 were referred to in section 4.2 of the main manuscript.
@@
Figure 9: Posterior coverage (top) calculated as the percentage of times the true parameter fell within the modeled 90% credible interval over the 500 replications. Coverage is faceted by the exponential parameter $\lambda$ and $n$ with $K=23$, and by increasing sample size ($x$-axis). The five models QGP, ORD, QGPN, ORDN, and IND are colored as shown the legend. The horizontal line (black) is at the nominal 90% level. Only QGP and ORD appear for the parameter $n$ as they are the only two which estimate an unknown $n$. (bottom) Distance between the true distribution and the estimated QM predictive distribution averaged over the 500 replications. Distances include UWD1, TV, and KLD for $K=23$, and by increasing sample size ($x$-axis).
Figure 10: Boxplots of time to fit a model in seconds ($y$-axis) for the 500 replicates in the simulation study for QM of the generalized lambda distribution (GLD) for $K\in\{3,7,15,23,50\}$ ($x$-axis). Boxplots are separated by QM methods QGP (white) and ORD (dark grey). The $y$-axis is on the $\text{log}_{10}$ scale.
@@
[…]

### v2

@@ 5 CDC flu forecasts analysis
In this section, the QGP and other models are used for QM of quantile forecasts targeting US flu hospitalizations. Beginning in 2013, the CDC began hosting a yearly forecast competition of the influenza outbreak in the US. This competition is known as FluSight. The flu epidemic typically begins in the fall and ends in late spring the following year, and the forecast competition lasts around 30 weeks starting in October and ending in May. FluSight involves a few dozen academic and industry research teams who each independently develop forecasts every week for predicting certain flu targets for future weeks [66]. During the 2022-23 and 2023-24 seasons, the forecast targets were the 1, 2, 3, and 4-week ahead hospitalizations as reported by Health and Human Services (HHS) for the 50 US states, Puerto Rico, the District of Columbia, and the nation as a whole. The data for weekly hospitalizations of flu patients may be found at [67]. The official guidelines of the 2023-24 FluSight and all submitted quantile forecasts are publicly available at [12] [11].
Participating teams were free to create forecasts however they pleased, but forecasts for each target were required to consist of 23 quantiles for probability levels $\boldsymbol{p}=(0.01,0.025,0.5,0.1,...,0.95,0.975,0.99)$, which were given by FluSight. Figure 6 shows examples of quantile forecasts of the same target from 12 participating teams. Each plot shows the log forecast for hospitalizations in the US for the week of January 13, 2024, and it is clear that there are major distributional differences between forecasts. We denote a quantile forecast for flu hospitalizations as $\hat{Q}^{(H)}(\boldsymbol{p})=(\hat{Q}^{(H)}(0.01),...,\hat{Q}^{(H)}(0.99))$. The general quantile forecast representation format allowed for forecasts to be compared by the same metric, the weighted interval score (WIS), and to be easily combined into a multi-model ensemble forecast [11]. However, under the quantile representation, the tools for scoring forecasts and for building ensemble models are limited as many of the existing tools for scoring forecasts and constructing ensembles require CDFs or PDFs [68, 69]. Because of these limitations, it may be desirable to approximate continuous distributions from the quantile forecasts to allow for more flexibility in scoring or ensemble building. In this section, we fit the QGP and other models for QM to forcasts submitted to the FluSight during the 2023-24 season, and an analysis is made to compare the fits.
@@
Before fitting a QGP model to a quantile forecast we transformed the quantiles $\hat{Q}^{(H)}(\boldsymbol{p})$ to $\text{log}(\hat{Q}^{(H)}(\boldsymbol{p})+1)$ so that forecasts for all states were on a similar scale. The model fit was the same as the QGP fit in section 4.2, including the same prior distribution assignments. According to the official forecast competition rules, no forecast could include quantiles that were less than 0. As a result, there were forecasts with one or many quantiles equal to 0. When fitting the QGP model, the 0 values were removed so that some forecasts had $K<23$ quantiles. Between 39 competing forecast teams, 53 locations, 29 forecast dates, and 4 horizons, there were 161,093 quantile forecasts on which QM was done. For each forecast, the WIS was calculated for the quantile forecast and the continuous ranked probability score (CRPS) was calculated over the predictive distribution of the fit QM models.
The WIS and CRPS are both proper scoring rules, which is a class of scoring rule defined so as to keep a forecaster honest in their forecasts [17, 18]. The definition for the WIS is in (19) and is the same as that in [16]. The WIS consists of the sum of multiple intervals scores (IS), the definition of which is in (20). Here $\alpha$ is the nominal level of an interval with $l$ and $r$ being the respective lower and upper bounds of the interval. $R$ is the number of intervals in the quantile forecast, $\alpha_{r}$ is the nominal level for the $r^{th}$ interval, and $y^{*}$ is the observed value which one is attempting to forecast.
@@
The CRPS is a widely used scoring rule for forecasting and is a function of the CDF of a continuous distribution thus making it unavailable for scoring quantile functions. The CRPS is defined in (21) where $F$ is the CDF of a forecast and $y^{*}$ is the observed event one attempts to forecast. The definition is the same as in [18].
$\text{CRPS}(F,y^{*})=\int_{-\infty}^{\infty}(F(x)-\mathds{1}(y^{*}\leq x))^{2}dx$ (21)
@@
Fitting continuous distributions to the quantile forecasts would allow for forecast comparison using the CRPS. The CRPS assesses a forecast across an entire distribution, including the tails, which the WIS is unable to do, giving more reason why one may want to perform QM on a quantile forecast. We fit the QGP model in (12) to the forecast competition forecasts from the 2023-24 season and compare the results of scoring the given quantile forecasts by the WIS and the CRPS calculated from the posterior predictive of the QGP model. Posterior predictive samples from the QM forecasts allow for approximate calculation of the CRPS. To calculate the WIS, we use the evalcast R package [70], and to calculate the CRPS we use the scoringutils package [71].
Figure 7 shows results of comparing the WIS from quantile forecasts and the CRPS from the fitted forecasts for QGP, ORD, IND, SPL, and KDE methods. The plots have in the $x$-axis the WIS and CRPS is in the $y$-axis. Each point represents one forecast for any participating team, season week, and state. Unsurprisingly, the correlation between the QM CRPS and WIS is very highly correlated for each method. For all but the KDE method, there is a tendency for some CRPS values to go above the $x=y$ line suggesting the CRPS scores are enlarged due to the addition of distribution tails. The CRPS scores for the IND method also go below the $x=y$ line suggesting that the method’s fit may be simply be poorer than the other methods. The slightly lower correlation for the QGP compared to ORD and IND may also be a result of adding more uncertainty in the fitting of tail quantiles.
@@
Figure 7: Scatterplots of WIS and CRPS values for forecasts from 39 different teams for 2023-24 CDC flu forecast competition. Each plot is for a different QM method. Each point represents scores for a flu forecast for one week during the season, one state, and one competing team. Points are transparent to show where more scores tended to be. Overall linear correlation is also given in the corner of each plot.
We do not know the true distributions from which each quantile forecast was made, but we can assess closeness of fit by measuring the difference between the QM quantiles and the original forecast quantiles given the FluSight probability values. Figure 8 shows the average over week mean absolute error (MAE) and mean sqaure error (MSE) between the the newly estimated QM quantiles and the original forecast quantiles. The MAE and MSE values for the SPL method are 0 for every week because the estimated quantile functions are interpolations of the original forecasts. Besides the SPL, however, the QGP appears to estimate quantiles that are the closest to the original forecast quantiles. The average MAE for the QGP is smaller than for the remaining methods for each week and the MSE is smaller for most weeks showing a robustness which the other methods – namely ORD and KDE – may lack.
Figure 8: Mean absolute error (MAE) and mean square error (MSE) over week for FluSight QM quantiles for each of QGP, ORD, IND, SPL, and KDE methods.
@@ 6 Conclusion
In this manuscript we considered the situation where quantiles with known probability levels are given as data or the quantiles represent a distribution rather than raw data or a fully defined distribution. We reviewed basic properties of QFs, sample quantiles, and resulting CLTs which motivate a new model, the QGP model, for matching a set of quantiles to a continuous distribution. The QGP model is assessed for parameter estimation and inference, distribution approximation, and it is compared with other QM methods found in the literature. The QGP provides explicit and easily interpretable uncertainty quantification of QM distributions, and generally outperforms the methods to which it is compared, sometimes with the exception of the ORD. An added advantage, however, is that the QGP provides a way to perform QM for distributions defined by CDFs as well as quantile distributions where the CDF is difficult to evaluate. An application of QM on quantile forecasts from the 2023-24 FluSight project demonstrates that QM methods allow for the use of the CRPS for scoring forecasts which is able to account for distribution tail uncertainty. QGP and other QM forecasts may be scored using a variety of scoring rules unavailable to quantile forecasts, and independent forecasts may be combined using methods made only for combining distributions by CDF or PDF functions. QGP has already been applied in chapter 4 of [72] who used the QGP predictive distributions fit in section 5 to construct ensemble forecasts. In the analysis, it is shown that the QGP may offer a more robust fit to quantile distributions than other QM methods.
Approximating continuous distributions given only a set of estimated quantiles is not a new idea, but QGP provides a method of doing so while also accurately accounting for asymptotic uncertainty in estimated or predicted quantiles. The CLTs in this manuscript are limited to the case where the quantiles are estimated given a distribution sample, however similar results have been shown for the quantile regression case where quantiles are estimated in the presence of covariates [47, 5]. These results could lead to additional QM modeling using the QGP model where covariates are given, and this may be more akin to the QM done by [27]. Additional research for the QGP model may be in how the required distribution function is selected. Where a true model is unknown, we elected to model a distribution as a normal mixture distribution. Of course this has its limits, and using other non parametric functions, similar to [39], may provide needed flexibility. As long as the selected function is continuous and once differentiable, the theory herein applies. The applications here suggests QM being useful for forecast hubs. [21] found QM a necessity in order to evaluate forecasts by a new scoring rule with COVID-19 specific applications.
Acknowledgments
@@
Data Availability
All data used in this manuscript may be found at https://github.com/cdcepi/FluSight-forecast-hub
All code used to obtain results in this manuscript, including code for accessing data, may be found at https://github.com/wadspen/quantile_fitting
@@
Supplementary Material
Additional simulation study results to section 4.1
Figure 9 shows examples of marginal posterior densities for QM of quantiles for a normal distribution from the simulation study in section 4.1 in the main manuscript. The parameter uncertainty of the QGP and ORD models are unsurprisingly similar in most cases, and where they differ it is hard to say that one is estimating the true parameter better than the other. The IND model on the other hand has tighter densities, and often the bulk of the posterior is far from the true parameter.
Figure 9: Density plots of posterior distribution samples for normal parameters by QM for QGP, ORD, and IND models. QM was done on estimated quantiles from a normal distribution with mean 4 and standard deviation 3.5. The posterior densities are for $\mu$ (top left), $\sigma$ (top right), and sample size n (bottom). Plots are faceted by the sample size $n\in\{50,150,500,1,000\}$ ($y$-axis) and number of quantiles $K\in\{7,15,23\}$ ($x$-axis). Vertical lines (black) show the value of the true parameter
Figures 10 and 11 are from a simulation study similar to that of section 4.1 except that QM was performed on quantiles estimated from draws from an exponential distribution with parameter $\lambda=4$ rather than a normal distribution. The results in the study for the exponential distribution were similar to those for the normal distribution.
@@
Figure 11: Posterior coverage (top) calculated as the percentage of times the true parameter fell within the modeled 90% credible interval over the 500 replications. Coverage is faceted by the exponential parameter $\lambda$ and $n$ with $K=23$, and by increasing sample size ($x$-axis). The five models QGP, ORD, QGPN, ORDN, and IND are colored as shown the legend. The horizontal line (black) is at the nominal 90% level. Only QGP and ORD appear for the parameter $n$ as they are the only two which estimate an unknown $n$. (bottom) Distance between the true distribution and the estimated QM predictive distribution averaged over the 500 replications. Distances include UWD1, TV, and KLD for $K=23$, and by increasing sample size (x-axis)
Additional simulation study results to section 4.2
Figures 12 and 13 show examples of QM fits to Laplace and normal mixture distributions as described in section 4.2.
@@
Figure 13: QM fits of $K=23$ quantiles by KDE, SPL, IND, and QGP for $n\in\{50,150,500,1{,}000\}$. The quantiles were sampled from the two component normal mixture distribution wN(1,0.9)+(1w)N(1.2,.6)wN(-1,0.9)+(1-w)N(1.2,.6) where $w=0.35$. The quantile fits (left) show the true quantiles (black) with either the QM fit line (grey) or the credible intervals of 50% (red) and 95% (pink). The estimated PDF plots (right) show the true PDF (black) with either a the QM estimated PDF (grey) or the credible intervals of 50% (red) and 95% (pink).
QGP normal mixture components analysis
This section shows an analysis which was made to determine the number of normal components to use in the distribution assigned for use in the QGP. Figure 14 shows the UWD1, TV, and KLD distances between a QGP fit distributions and simulated quantiles averaged over 500 simulated replicates. Table 1 shows the average values. The results show that while the larger the number of components, the closer is the fit distribution in terms of the given distances. However, the imrpovement after three or four components is not much making four components attractive for relatively good fit and for model simplicity.
Figure 14: UWD1, TV, and KLD distances averaged over 500 simulation replicates for QGP fits of quantile simulated from extreme value, Laplace, and mixture distributions. The QGP with normal mixture distributions of 1 to 6 components were fit for both the cases where $n$ is not known (left) and $n$ is known.
Table 1: UWD1, TV, and KLD distances averaged over 500 simulation replicates for QGP fits of quantile simulated from extreme value, Laplace, and mixture distributions. The QGP with normal mixture distributions of 1 to 6 components were fit for both the cases where $n$ is not known (left) and $n$ is known.
QGP QGP-n Components 1 2 3 4 5 6 1 2 3 4 5 6 UWD1 EV 0.063 0.025 0.021 0.020 0.020 0.020 0.064 0.025 0.021 0.021 0.020 0.020 La 0.044 0.022 0.020 0.020 0.020 0.020 0.044 0.021 0.020 0.020 0.020 0.020 MIX 0.107 0.018 0.019 0.019 0.019 0.019 0.157 0.018 0.019 0.018 0.018 0.018 TV EV 0.127 0.053 0.047 0.047 0.048 0.049 0.125 0.052 0.039 0.039 0.039 0.039 La 0.117 0.051 0.047 0.047 0.048 0.048 0.117 0.048 0.043 0.043 0.043 0.042 MIX 0.226 0.038 0.033 0.032 0.033 0.034 0.225 0.029 0.031 0.032 0.034 0.035 KLD EV 0.171 0.025 0.014 0.013 0.014 0.015 0.170 0.025 0.011 0.010 0.010 0.010 La 0.110 0.014 0.012 0.012 0.012 0.012 0.111 0.013 0.010 0.010 0.010 0.010 MIX 0.221 0.010 0.008 0.008 0.009 0.010 0.293 0.004 0.006 0.007 0.009 0.010